Mole Problems

  1. Calculate the mass of 2.5 moles of NaCl.
  2. Determine the mass in grams of 0.500 moles of Ca(OH)2.
  3. Calculate the mass of 0.010 moles of NH3.
  4. Determine the mass of 3.0 moles of CO2.
  5. Calculate the number of moles in 168.0 g of HgS.
  6. Determine the number of moles in 27.0 g of H2O.
  7. Calculate the number of moles in 50.0 g of NH3.
  8. Calculate the mass of 6.022 x 1023 particles of CaSO4.
  9. Convert 9.500 x 1024 molecules of CH4 into a mass in grams.
  10. Determine the mass of 3.011 x 1023 atoms of gold.
  11. Calculate the number of molecules in 15.0 grams of CaSO4.
  12. Determine the number of molecules in 100.0 g of KNO3.
  13. Which would have more molecules, 2.0 moles of hydrogen gas, H2, or 2.0 moles of methane gas, CH4? Explain your answer.
  14. Which would have the larger mass, 2.0 moles of hydrogen gas, H2, or 2.0 moles of methane gas, CH4? Explain your answer.
1
Expert's answer
2021-03-23T07:25:09-0400

1. m(NaCl) = M×n

m(NaCl) = 58.5×2.5 = 146.25 g

2. m((Ca(OH)2) = 0.500×74 = 37 g

3. m(NH3) = 0.010×17 = 0.17 g

4. m(CO2) = 3.0×44 = 132 g

5. n(HgS) = m/M

n(HgS) = 168.0/233 = 0.72 mole

6. n(H2O) = 27.0/18 = 1.5 mole

7. n(NH3) = 50.0/17 = 2.94 mole

8. n(CaSO4) = N/NA

n(CaSO4) = 6.022×1023/6.022×1023 = 1 mole

m(CaSO4) = 136 g

9. n(CH4) = 9.500×1024/6.022×1023 = 15.77 mole

m(CH4) = 15.77×18 = 284 g

10. n(Au) = 3.011×1023/6.022×1023 = 0.5 mole

m(Au) = 0.5×197 = 98.5 g

11. n(CaSO4) = m/M

n(CaSO4) = 15.0/136 = 0.11 mole

N = n×NA

N = 0.11× 6.022×1023 = 0.66×1023 molecules

12. n(KNO3) = 100/101 = 0.99 mole

N(KNO3) = 0.99 × 6.022×1023 = 5.96×1023 molecules

13. The number of molecules in a mole is constant and is equal to Avogadro's number: 6.022×1023.

Therefore, 2.0 mole of H2 and 2.0 mole CH4 all contain the same number of molecules which is 6.022×1023 molecules.

14. m(H2) = 2.0×2 = 4 g

m(CH4) = 2.0×18 = 36 g

CH4 have the larger mass.